# Improving Referring Image Segmentation using Vision-Aware Text Features

### v1

@@ 3.4 Network Training
Training. Our prediction $\hat{y}_{\delta}$ is supervised by three losses. Firstly, the class loss $\mathcal{L}_{cls}$ is binary cross entropy (BCE) loss to supervise the probability of referred object. Secondly, the mask loss $\mathcal{L}_{mask}$ is a combination of dice loss and BCE. Finally, our sentence-level contrastive loss $\mathcal{L}_{mcc}$ is used to enforce our Meaning Consistency Constraint. The total loss can be formulated as follows:
@@ 4.1 Experiment Setup
Method Backbone RefCOCO RefCOCO+ G-Ref
Visual Textual val testA testB val testA testB val test
VLT [8] Darknet-53 Bi-GRU 65.65 68.29 62.73 55.50 59.20 49.36 52.09 56.65
LAVT [52] Swin-B BERT 74.46 76.89 70.94 65.81 70.97 59.23 63.34 63.62
JMCELN [17] ResNet-101 CLIP 74.40 77.69 70.43 66.99 72.69 57.34 64.08 64.99
VG-LAW [40] ViT-B BERT 75.05 77.36 71.69 66.61 70.30 58.14 65.36 65.13
We evaluate the performance of our model on three image datasets: RefCOCO [18], RefCOCO+ [18], G-Ref [36] and further evaluate the performance of our model on two video datasets: Ref-Youtube-VOS [39] and Ref-DAVIS17 [19].
For evaluation metrics, we calculate mIoU to measure the overlap between the predicted segmentation mask and the ground truth mask for image segmentation. We also calculate region similarity $\mathcal{J}$ and contour accuracy $\mathcal{F}$, and their average $\mathcal{J}\&\mathcal{F}$ for video object segmentation.
We use PyTorch to implement our method. During training, we freeze both the CLIP Visual and Text Encoder. Images are resized to a short side of 480. We set the coefficients for the losses as $\gamma_{cls}=2,\gamma_{mask}=5$, and $\gamma_{mcc}=2$, with the feature dimension $C$ set to $256$. We train the network for $100,000$ iterations on the RefCOCO(+/g) datasets with an initial learning rate of $10^{-4}$ and is reduced by a factor of $0.1$ at the $2/3$ last iteration. For the Ref-Youtube-VOS dataset, we initialize the pre-trained weight from RefCOCO(+/g) and train the network for $100,000$ iterations. Regarding the Ref-DAVIS17 dataset, we directly use the weight obtained from the Ref-Youtube-VOS dataset for inference. For detailed information on each dataset and implementation, please see the supplementary material.
Table 2: Quantitative results on video dataset using Video-Swin-B backbone.
Methods Ref-YT-VOS Ref-DAVIS17
[…]
ReferFormer [46] 62.9 61.3 67.5 61.1 58.1 64.1
VLT  [9] 63.8 61.9 65.6 61.6 58.9 64.3
VATEX (Ours) 65.4 63.3 67.5 65.4 62.3 68.5
Table 3: Performance comparison between LAVT and VATEX on Oc-IoU metric.
Method RefCOCO RefCOCO+ G-Ref
LAVT [52] 62.51 50.79 56.01
Ours w/o MCC 66.42 54.92 59.25
Ours 68.20 57.38 61.69
4.2 Main Results
Referring Image Segmentation. We compare our proposed method with the state-of-the-art methods on three common datasets (RefCOCO, RefCOCO+, G-Ref). As illustrated in Table 1, our method outperforms the state-of-the-art methods by a large margin in all splits of different datasets in the standard setting. Particularly, for the RefCOCO dataset, our method outperforms the recent VG-LAW, with mIoU gains of 3.11%, 2.28%, and 3.95% across the three splits, respectively. For the more complex RefCOCO+ dataset, our method outperforms the previous top method, JMCELN, by mIoU margins of 3.03%, 1.72%, and 5.18% in the validation, test A, and test B sets, respectively. The G-Ref dataset presented the most significant challenge, with longer and more complicated language expressions. Despite this, our method achieved remarkable performance improvements of approximately 5% in mIoU on the validation and test sets, compared to the second-best method, VG-LAW, in the same setting.
Referring Video Segmentation. Our model can be extended to video datasets with minor adaptations to handle temporal information. As shown in Table 3, VATEX outperforms the current SOTA methods VLT and ReferFormer on the same Video-Swin-B backbone by 1.6 and 3.8 $\mathcal{J}\&\mathcal{F}$ on the Ref-Youtube-VOS and Ref-DAVIS17 datasets, respectively.
@@ 4.2 Main Results
Table 5: Analysis of prompt templates in CLIP Prior on RefCOCO validation set.
Query Initialization method RefCOCO (mIoU)
Text features as query [46] 75.43
CLIP Prior (Naive Implementation) 74.34 -1.09
@@
78.25 +2.82
@@ 4.3 More Analysis
CLIP Prior Analysis. As described in Section 3.1, our CLIP Prior relies on a template sentence to convert a complex sentence into a simplified sentence suitable for CLIP. The baseline which uses text features as initial query achieves 75.43% mIoU. We investigate the effects of using different templates and how these affect the final performance. Table 5 demonstrates that using the complex sentence leads to a significant deterioration in performance at 1.09%. In this case, CLIP introduces noisy activation on various objects based on their discriminative characteristics within the complex sentence, which is harmful to the model. By using the template “A Photo of [Object]”, there is a notable improvement of 2.73% in mIoU. We conducted an additional experiment where we aggregated the text embeddings from 80 ImageNet prompts, which has a very minor performance improvement. This demonstrates that the template “A Photo of [Object]” remains a practical choice.
Regarding the quality of the heatmap, Figure 8 demonstrates the enhanced localization performance achieved when prompt templates are utilized, compared to the original expression. Without simplifying the sentence, the heatmap tends to encompass distinctive categories in images or mentioned in the sentence (e.g. obstacles, woman and blue shirts), or highlights the wrong object (e.g. wind-surfing boat). By reducing the complexity of the text expression, it can be seen that the activation of the object of interest becomes more accurate. For more visualization, please see the supplementary material.
@@
Figure 6: Our heatmap from CLIP Prior.
Effect of CMD and MCC on Vision-aware Text Features. To study the impact of the CMD and MCC on text features, we calculate the similarity between sentence-level text features before and after processing through CMD. The average similarity results for all pairs of expressions referring to the same or different objects at each CMD level are depicted in Figure 6. Here, Level 0 represents the initial features, derived directly from the CLIP Embedding, while Level 4 indicates the final vision-aware text features of CMD.
In the context of expressions referring to the same object, the initial similarity of text features stands at $68.88\%$, $65.82\%$, and $63.35\%$ for RefCOCO, RefCOCO+, and G-Ref, respectively. Following just one CMD level, the similarities witness an approximate $7\%$ elevation across all datasets. The final vision-aware text features of the CMD process yield similarities of $81.29\%$ in RefCOCO, $78.28\%$ in RefCOCO+, and $76.19\%$ in G-Ref.
Regarding different objects, we compute the similarity between all expressions pertaining to different objects within the same image to showcase the influence of our CMD module on them. As illustrated in Figure 6(b), the similarity between the features of these expressions progressively decreases from $53.23\%$, $55.23\%$, $60.06\%$ to $22.12\%$, $32.82\%$, $36.81\%$ on RefCOCO, RefCOCO+, and G-Ref, respectively.
These findings underscore the pivotal role of CMD in bolstering multimodal comprehension. Specifically, we reveal how CMD effectively improves the representation of text features that are conditioned on visual context, illustrating its significance in advancing the interaction between textual and visual information.
Effect of MCC on Object segmentation mask. To quantify the ability to consistently segment various expressions for the same object and further validate the effectiveness of our proposed Meaning Consistency Constraint, we leverage an Object-centric Intersection over Union (Oc-IoU) score, which calculates the overlap and union area between ground truth and all segmentation predictions of the same object. Specifically, consider the $i$-th object with $K_{i}$ expressions referring to that object and the corresponding ground truth mask $\text{GT}_{i}$. Let $P_{i}^{j}$ be the model’s prediction for the $j$-th expression of the $i$-th object, where $j=\overline{1..\;K_{i}}.$ The Object-centric IoU can be formulated as follows:
@@
[…]
where $N$ is the total number of objects/instances in the datasets.
Table 3 provides the comparisons between our method and the state-of-the-art method LAVT in Oc-IoU on the validation set of three RIS benchmarks. As can be seen, our method outperforms LAVT in all three datasets. Comparing the last two rows of Table 3, we can see that the MCC helps the model, especially CMD to enhance mutual information between textual and visual features to further provide more consistent and accurate segmentation. These results underscore the compelling efficacy of our Meaning Consistency Constraint in resolving language ambiguities, thus improving the segmentation performance.
Qualitative Analysis. We provide the visualization of our results in Figure 7. Our method can successfully segment objects in complex scenarios, such as the presence of multiple similar objects. For example, in the first column, we can localize the guy who sits in the chair instead of the man standing on the tennis court. In the second sample, VATEX can distinguish the sushi plate among several food dishes that LAVT cannot. In the fourth column, our model can not only identify the correct umbrella belonging to the guy in the light jacket but also segment a part of the shaft of the umbrella that the ground truth does not provide. With the expression "bear child is hugging" in the sixth image, LAVT can only segment the bear’s head, but VATEX can output the whole annotation of the target bear. However, our method fails to segment objects that need to be counted and selected by their order or be described indirectly through another object, as we have not leveraged counting information and object interaction in our model.
Another point worth mentioning is the differences in architecture design between VATEX and LAVT, with VATEX focusing on instance-based segmentation and LAVT on pixel-based segmentation. This will lead to smoother, more complete segmentation masks by VATEX. However, this strength becomes a weakness when the model misidentifies an instance, resulting in entirely incorrect segmentation masks, whereas LAVT’s pixel-based approach can still partially identify the object in such situations.
Figure 7: Results on RefCOCO(+/g) datasets. We compare our results with LAVT. Our method excels at segmenting objects in complex scenarios, such as distinguishing between similar objects and localizing specific instances within a scene. The last two columns of the results show failure cases. Best viewed in color.
5 Limitations and Future Works
Our method is not without limitations. Particularly, our method does not exploit the interaction and relationship between different objects as well as the alignment between actions and expressions referring to them (see Figure 7 the last two columns). Consequently, situations involving counting ("third from left"), indirect descriptions ("kid next to girl in pink pants"), or actions ("a woman walking to the left") might lead to inaccurate predictions. Dealing with intra-frame object relationships and inter-frame information is vital for future work. Another line of future work is making RIS work on general scenarios (e.g. segment all the red-colored objects, segment all the text in the image) or more fine-grained segmentation (e.g. segment the eye of the owl). It is also of great interest to investigate vision-aware text features with the SOTA vision-and-language model and to lift the referring expression segmentation task to the 3D domain.

### v2

@@ 3.4 Network Training
Training. Our prediction $\hat{y}_{\delta}$ is supervised by three losses. Firstly, the class loss $\mathcal{L}_{cls}$ is binary cross entropy (BCE) loss to supervise the probability of the referred object. Secondly, the mask loss $\mathcal{L}_{mask}$ is a combination of dice loss and BCE. Finally, our sentence-level contrastive loss $\mathcal{L}_{mcc}$ is used to enforce our Meaning Consistency Constraint. The total loss can be formulated as follows:
@@ 4.1 Experiment Setup
We evaluate the performance of our model on three image datasets: RefCOCO [24], RefCOCO+ [24], G-Ref [42] and further evaluate the performance of our model on two video datasets: Ref-Youtube-VOS [45] and Ref-DAVIS17 [25]. For evaluation metrics, we follow previous work to use mean IoU (mIoU) and Precision at different thresholds (Pr@X) for image and $\mathcal{J}\&\mathcal{F}$ for video datasets.
During training, we freeze the CLIP model. Images are resized to a short side of 480. We set the coefficients for the losses as $\gamma_{cls}=2,\gamma_{mask}=5$, and $\gamma_{mcc}=2$, with the feature dimension $C$ set to $256$. We train the network for $100,000$ iterations on the RefCOCO(+/g) datasets with an initial learning rate of $10^{-4}$ and is reduced by a factor of $0.1$ at the $2/3$ last iteration. For the Ref-Youtube-VOS dataset, we initialize the pre-trained weight from RefCOCO(+/g) and train the network for $100,000$ iterations. Regarding the Ref-DAVIS17 dataset, we directly use the weight obtained from the Ref-Youtube-VOS dataset for inference. The training process uses 2 NVIDIA RTX 3090 GPUs with a batch size of 24. For detailed information on each dataset and implementation, please see the supplementary material.
Figure 5: Results on RefCOCO(+/g) datasets. We compare our results with CRIS and LAVT. Our method excels at segmenting objects in complex scenarios, such as distinguishing between similar objects and localizing specific instances within a scene. The last two columns of the results show failure cases. Best viewed in color.
@@
Table 5: Ablation on different query initialization methods in CLIP Prior.
Query Initialization method RefCOCO
Only text features [54] 75.43 -2.73
@@
78.25 +0.09
CLIP Prior w.o main object extractor 74.34 -3.82
Table 6: Ablation on different bidirectional multimodal fusion modules.
Bidirectional fusion MCC RefCOCO
CMD (Ours) 78.16
ETRIS [58] 77.22 -0.94
CoupAlign [64] 77.01 -1.15
CMD 75.12 -3.04
ETRIS [58] 74.12 -4.04
CoupAlign [64] 73.97 -4.19
4.2 Main Results
Referring Image Segmentation. As illustrated in Table 3, our method outperforms the state-of-the-art methods by a large margin in all splits of different datasets in the standard setting. Notably, our method surpasses the recent CGFormer and VG-LAW on the validation splits of all three benchmarks, achieving mIoU gains of 1.23% and 3.11% on RefCOCO, 1.46% and 3.31% on RefCOCO+, and 2.16% and 4.37% on G-Ref. The more complex the expressions, the greater the performance gains achieved by VATEX. Compared to LISA [29], a large pre-trained vision and text encoder, VATEX consistently outperforms it by 3-5% across all datasets. Furthermore,  Table 3 demonstrates the superior performance of VATEX over LAVT, ReLA, and CG-Former on average precision metrics, particularly at the Pr@0.7 and Pr@0.9, illustrating our ability to generate high-quality and complete segmentation masks.
Referring Video Segmentation. Our model can be extended to video datasets with minor adaptations to handle temporal information. As shown in Table 3, VATEX outperforms current SOTA methods VLT and ReferFormer on the same Video-Swin-B backbone by 1.6 and 3.8 $\mathcal{J}\&\mathcal{F}$ on Ref-Youtube-VOS and Ref-DAVIS17 datasets, respectively.
Qualitative Analysis. We provide the visualization of our results in Figure 5. Our method can successfully segment objects in complex scenarios, such as the presence of multiple similar objects. For example, in the first column, we can localize the guy who sits in the chair instead of the man standing on the tennis court. In the second sample, VATEX can distinguish the sushi plate among several food dishes that LAVT cannot. In the fourth column, our model can not only identify the correct umbrella belonging to the guy in the light jacket but also segment a part of the shaft of the umbrella that the ground truth does not provide. With the expression ”bear child is hugging” in the sixth image, LAVT can only segment the bear’s head, and CRIS over-segment to the child and under-segment the bear, but VATEX can output the target bear concisely. However, our method fails to segment objects that need to be counted and selected by their order or be described indirectly through another object, as we have not leveraged counting information and object interaction in our model. Another point worth mentioning is the differences in architecture design between VATEX and LAVT. VATEX focuses on instance-based segmentation, while LAVT focuses on pixel-based segmentation. Consequently, VATEX produces smoother and more complete segmentation masks.
@@ 4.3 More Analysis
Study on different query initialization methods of CLIP Prior. As described in Section 3.1, our CLIP Prior relies on a template to convert the main noun phrase into a simple sentence suitable for CLIP. As shown in Table 6, the baseline follows the query initialization from  [54], which uses only text features and achieves 75.43% mIoU. We investigate the effects of using different templates and how these affect the final performance. By using the template “A Photo of [Object]”, there is a notable improvement of 2.73% in mIoU. We conducted an additional experiment where we aggregated the text embeddings from 80 ImageNet prompts, which has a very minor performance improvement. This demonstrates that the template “A Photo of [Object]” remains a practical choice. We also conduct the experiment that using the whole sentence (instead of the main noun phrase) leads to a significant deterioration in performance at 3.82%. In this case, CLIP introduces noisy activation on various objects based on their discriminative characteristics within the complex sentence, which is harmful to the model.
Different bidirectional multimodal fusion modules. We ran an ablation study to quantify the performance of CMD with other bidirectional multimodal fusion mechanisms by substituting CMD with alternative modules in ETRIS [58] and CoupAlign [64]. As shown in Table 6, CMD with MCC performs the best, achieving a mIoU of $78.16\%$, which is higher than both ETRIS [58] and CoupAlign [64] with $0.94\%$ and $1.15\%$, respectively. Disabling MCC results in a notable drop in performance across all settings, with CMD seeing a decrease of $3.04\%$, while ETRIS and CoupAlign experience decreases of $4.04\%$ and $4.19\%$, respectively. This highlights the importance of MCC in improving bidirectional fusion performance, with CMD consistently outperforming alternative methods under the same condition.
Effect of MCC on Vision-aware Text Features. To quantify the impact of the MCC on Vision-aware Text Features, we calculate the similarity between sentence-level text features at each layer of CMD, with and without MCC, using the G-Ref dataset. We chose G-Ref because it contains longer, more diverse, and complex context information about the objects expressions, making it ideal for studying the impact of MCC on Vision-aware Text Features. The average similarity results for all pairs of expressions referring to the same or different objects are depicted in Figure 6. Here, level 0 represents the initial features, derived directly from the text encoder, while level 4 indicates the final vision-aware text features of CMD.
@@
Regarding different objects, the pairwise similarity of expressions referring to different objects progressively decreases through the four levels of CMD, from 60.06% to 36.81%. In contrast, using CMD without MCC results in an unstable feature space. This illustrates the effectiveness of MCC in guiding vision-aware text features toward a more robust feature space, where the features of two sentences referring to different objects can be similar at the beginning but are pushed apart in the final vision-aware text feature.
These findings underscore the pivotal role of MCC in bolstering multimodal comprehension provided by CMD. Specifically, we reveal how the in-context learning signal in MCC effectively improves the representation of vision-aware text features.
Others. We reported the Universality of VATEX, the Runtime Analysis of VATEX, and the effect of MCC on object segmentation in Supplementary Material.
5 Limitations
Our method is not without limitations. Particularly, our method does not exploit positional relationship between different objects as well as the alignment between actions and expressions referring to them (see Figure 5 the last two columns). Consequently, situations involving counting (”third from left”), indirect descriptions (”kid next to girl in pink pants”), or actions (”a woman walking to the left”) might lead to inaccurate predictions. Another line of future work is making RIS work on general scenarios (e.g. segment all the red-colored objects, segment all the text in the image) or more fine-grained segmentation (e.g. segment the eye of the owl). Dealing with intra-frame object relationships and inter-frame information for video is vital for future work. It is also of great interest to investigate vision-aware text features with the VLMs and to lift this task to the 3D domain.